OpenStudy (anonymous):

If my problem is -5r^2-40r-82 and its asking me to graph my equation but my graph only goes up to 10 and my vertex is (-8,-82) how do i do this

OpenStudy (anonymous):

Your vertex is incorrect. The x-value of your vertex can be found by leting x=-b/2a where b=-40 and a= -5. So x=-(-40/2(-5))=40/(-10)=-4. Then replacing x=-4 into your equation, y=-5(-4)^2-40(-4)-82=-80+160-82=-2. Thus, your vertex is well within your graph values. V=(-4, -2).